theronhoward400:

Given the function h(x) = 2(5)x, Section A is from x = 0 to x = 1 and Section B is from x = 2 to x = 3. Part A: Find the average rate of change of each section. (4 points) Part B: How many times greater is the average rate of change of Section B than Section A? Explain why one rate of change is greater than the other. (6 points) (10 points)

crispyrat:

@angle

Angle:

Hello @theronhoward400 ! Did you have a picture for your problem? is the function h(x) = 2 (5)^x ?

theronhoward400:

yes

theronhoward400:

no I don't

Angle:

so we plug in x = 0 and solve for y then we plug in x = 1 and solve for y and we are looking for what is essentially the slope between (0, y) and (1, y)

Angle:

y = 2 * 5^(0) = ? and y = 2 * 5^(1) = ?